how do you use zero product property to solve an equation like: (x-3)(x-2)=0?
Found 2 solutions by kiru_khandelwal, longjonsilver:
Answer by kiru_khandelwal(79)

For an equation like (x-3)(x-2) = 0
Zero product property means that the value of x shud be such that either (x-3) = 0 or (x-2) = 0, because then only the product will be 0
x-3 will be zero when x=3
and x-2 will be zero when x=2
therefore, (x-3)(x-2) = 0 has solution 3,2

Answer by longjonsilver(2297)

Give me 2 numbers that multiply to give zero.

2x3? nope
4x12? nope
-5x-56? nope

well?

how about 4x0? YES
or 27x0? YES
or -78x0? YES

To give zero when multiplied, one of the 2 numbers MUST be zero... so (x-3)(x-2)=0. One of (x-3) or (x-2) must be zero. Which one? Could be either, so we have to say that:

either x-3=0 or x-2=0

in which case, x=3 or x=2...ANSWER!

Jon.